4.​Suppose the demand for inject printers is estimated to be 

   Q= 1000-2p+5Px-3Py+0.1Y

  P=80   Px= 20     Py= 150       Y= 1000

A.​Calculate price elasticity of demand? (2 pt) 

B.​Calculate cross price elasticity of x and y (Exy)? And state the nature of good 

C.​Calculate income elasticity of demand and what we can say about the goods


1
Expert's answer
2021-07-12T16:17:52-0400

a) "Q= 1000- 2P+ 5PX- 3PY+ 0.1Y"

"Q= 1000-2(80) + 5(20) - 3(150) + 0.1(1000) = 590"

Price elasticity = change in Q/change in P*(P/Q)

"Ped=-2*80\/590 = -0.27"

b) Cross price elasticity of x= (change in Q/ Change in PX)* (PX/Q)

"5*20\/590 = 0.1695"

The good is a substitute since the cross-price elasticity is positive.

Cross price elasticity of y= (change in Q/ change in PY)* (PY/Q)

"-3*150\/590=- 0.7627"

The good is a compliment since its cross-price elasticity is negative.

c) Income elasticity of demand= (change in Q/ change in Y)*(Y/Q)

"0.1*1000\/590= 0.1695"

The goods are normal goods since their income elasticity of demand is positive.